40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 6 01:09:11 2017, johannes, Update, Computer Scripts / Programs, New set of loss measurements XARM_20170905.pdfYARM_20170905.pdf
    Reply  Mon Sep 11 12:56:40 2017, johannes, Update, Computer Scripts / Programs, lossmap attempts 
       Reply  Mon Sep 11 16:46:09 2017, Steve, Update, PEM, earthquakes M5.4eq.png
          Reply  Tue Sep 19 14:00:41 2017, Steve, Update, PEM, earthquakes 7.1MX5.3J.png
             Reply  Fri Oct 27 09:34:20 2017, Steve, Update, PEM, earthquakes recentEQs.png
                Reply  Fri Nov 10 13:46:26 2017, Steve, Update, PEM, M3.1 local earthquake 3.1M_local_EQ.png
                   Reply  Mon Apr 9 08:39:39 2018, Steve, Update, PEM, M5.3 eq Souther CA M5.3_Santa_Cruz_Is.CA.pngafter_M5.3.pngM5.3vac.png
                Reply  Thu Jan 25 08:27:25 2018, Steve, Update, PEM, earthquakes M4_local_eq.pngM4.pngM5.8Ferndale_CA.png
       Reply  Mon Sep 25 17:56:33 2017, johannes, Update, Computer Scripts / Programs, transmitted power during lossmap 
    Reply  Thu Oct 11 13:40:51 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
       Reply  Fri Oct 12 12:29:34 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
          Reply  Fri Oct 12 20:20:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
             Reply  Sat Oct 13 20:11:10 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                Reply  Mon Oct 15 10:32:13 2018, yuki, Update, Computer Scripts / Programs, loss measurements XARM_20181015_1500.pdfYARM_20181015_1500.pdf
                   Reply  Tue Oct 16 00:44:29 2018, yuki, Update, Computer Scripts / Programs, loss measurements 
                   Reply  Fri Nov 2 19:25:16 2018, gautam, Update, Computer Scripts / Programs, loss measurements 
    Reply  Mon Oct 15 12:52:54 2018, yuki, Update, Computer Scripts / Programs, additional comments 
       Reply  Fri Nov 2 16:42:31 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
          Reply  Mon Nov 5 13:52:18 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
             Reply  Tue Nov 6 10:19:26 2018, aaron, Update, Computer Scripts / Programs, arm loss measuremenents 
                Reply  Tue Nov 6 15:23:48 2018, gautam, Update, IOO, IMC problematic MCwonky.png
                   Reply  Tue Nov 6 19:02:35 2018, aaron, Update, IOO, IMC problematic 181106_MCTRANS.jpg
                      Reply  Sat Nov 10 17:40:00 2018, aaron, Update, IOO, IMC problematic 
                         Reply  Mon Nov 12 13:53:20 2018, rana, Update, IOO, loss measurement: oscope vs CDS DAQ 
                         Reply  Thu Nov 15 10:21:07 2018, aaron, Update, IOO, IMC problematic 181115_chansDown.pngPD_noise.png
                            Reply  Fri Nov 16 10:53:07 2018, aaron, Update, IOO, IMC problematic Screenshot_from_2018-11-16_19-22-34.png
                               Reply  Sat Nov 17 18:59:01 2018, aaron, Update, IOO, IMC problematic 
                Reply  Wed Nov 7 05:16:16 2018, yuki, Update, Computer Scripts / Programs, arm loss measuremenents 
Message ID: 14297     Entry time: Thu Nov 15 10:21:07 2018     In reply to: 14289     Reply to this: 14300
Author: aaron 
Type: Update 
Category: IOO 
Subject: IMC problematic 

I ran a BNC from the PD on the AS table along the cable rack to a free ADC channel on the LSC whitening board. I lay the BNC on top of the other cables in the rack, so as not to disturb anything. I also was careful not to touch the other cables on the LSC whitening board when I plugged in my BNC. The PD now reads out to... a mystery channel. The mystery channel goes then to c1lsc ADC0 channels 9-16 (since the BNC goes to input 8, it should be #16). To find the channel, I opened the c1lsc model and found that adc0 channel 15 (0-indexed in the model) goes to a terminator.

Rather than mess with the LSC model, Gautam freed up C1:ALS-BEATY_FINE_I, and I'm reading out the AS signal there.

I misaligned the x-arm then re-installed the AS PO PD, using the scope to center the beam then connecting it to the BNC to (first the mystery channel, then BEATY). I turned off all the lights.

I went to misalign the x-arms, but the some of the control channels are white boxed. The only working screen is on pianosa.

The noise on the AS signal is much larger than that on the MC trans signal, and the DC difference for misaligned vs locked states is much less than the RMS (spectrum attached); the coherence between MC trans and AS is low. However, after estimating that for ~30ppm the locked vs misaligned states should only be ~0.3-0.4% different, and double checking that we are well above ADC and dark noise (blocked the beam, took another spectrum) and not saturating the PD, these observations started to make more sense.

To make the measurement in cds, I also made the following changes to a copy opf Johannes' assess_armloss_refl.py that I placed in /opt/rtcds/caltech/c1/scripts/lossmap_scripts/armloss_cds/   :

  • function now takes as argument the number of averages, averaging time, channel of the AS PD, and YARM|XARM|DARK.
  • made the data save to my directory, in /users/aaron/40m/data/armloss/

I started taking a measurement, but quickly realized that the mode cleaner has been locked to a higher order mode for about an hour, so I spend some time moving the MC. It would repeatedly lock on the 00 mode, but the alignment must be bad because the transmission fluctuates between 300 and 1400, and the lock only lasts about 5 minutes.

Attachment 1: 181115_chansDown.png  52 kB  | Hide | Hide all | Show all
181115_chansDown.png
Attachment 2: PD_noise.png  49 kB  Uploaded Thu Nov 15 15:39:38 2018  | Show | Hide all | Show all
ELOG V3.1.3-